OpenStudy (anonymous):

taking my AP Calc class was a big mistake. i hate calculus and im usually a math wizard. I DO NOT understand the chain rule at all.... When do i use it? how do i use it? ESPECIALLY with trig functions.. do i add or multiply the derivatives? example 1) find the derivative (3x^2)(secx) example 2) find the derivative (3xsinx)+(x^2)(cosx) please tell me why you did what you did on each step. i seriously do not get this at all i seriously need help. i mean i got a 100% on my pre calc exam! why is this so hard for me :( please help thanks!

OpenStudy (anonymous):

first one use product rule, power rule for the exponents. There is no chain rule there

OpenStudy (anonymous):

Math wizard?

OpenStudy (anonymous):

\[ (fg)'=f'g+fg' \]

OpenStudy (anonymous):

1) 6xsecx+3x^2tanxsecx

OpenStudy (anonymous):

the second one also doesnt require the chain rule

OpenStudy (anonymous):

thank you azbits and wio. dear azbits, why doesnt the second example require a chain rule? how do i know when to use the chain rule?

OpenStudy (anonymous):

first of all, the chain rule. \[\frac{ d }{ dx }(f\circ g)(x) = f(g(x)) = f'(g(x))\cdot g'(x)\]

OpenStudy (anonymous):

2) 3(sinx+xcosx)+(2xcosx-x^2sinx) all done by product rule and power rule

OpenStudy (anonymous):

dear pgpiolet326, is that the basic chain rule formula? because i see stuff with a u' and that throws me off!

OpenStudy (anonymous):

dear azbits, how do i do the second example?

OpenStudy (anonymous):

you use the chain rule when you need to derive a function inside a function like f'(g(x)) examples sin(3x^2)+x^2(cos(3x^2)

OpenStudy (anonymous):

its so confusing :(

OpenStudy (anonymous):

dear azbits, ohhhh ok so we use the chain rule to avoid crazy steps? what about when i see a square root? when do i use the chain rule for that?

OpenStudy (anonymous):

sometimes books, teachers will write u for f(x) and v for g(x). this is especially true for the product rule.

OpenStudy (anonymous):

a simple example is sin(4x) if you derive that you get 4cos(4x)

OpenStudy (anonymous):

same square root = ^1/2

OpenStudy (anonymous):

dear pgpilot326, ohhh thats what throws me off thats so stupid when they add the u' :/

OpenStudy (anonymous):

ok thank you guys! and how do i know when to use the chain rule with the square root?

OpenStudy (anonymous):

savi if you have a square root like (x^2+1)^(1/2) then yes you can use the chain rule

OpenStudy (anonymous):

buts if its like sqrt(x) then you can just use the power rule

OpenStudy (anonymous):

dear azbits, is it because of the exponent? or should i use the chain rule everytime i see a square root? lol

OpenStudy (anonymous):

\[\frac{ d }{ dx }\sqrt{5x-6}=\frac{ d }{ dx }(5x-6)^{\frac{ 1 }{ 2 }}=\frac{ 1 }{ 2 }(5x-6)^{-\frac{ 1 }{ 2 }}\cdot(5)\]

OpenStudy (anonymous):

you should always convert roots to exponents

OpenStudy (anonymous):

for derivatives

OpenStudy (anonymous):

dear pgpilot326, ok thanks for the example. but do i ALWAYS use the chain rule with a square root?

OpenStudy (anonymous):

no you should only use chain rule when you have a functions that have more than one variable or when the variable doesnt appear by itself

OpenStudy (anonymous):

dear azbits, what does that mean?

OpenStudy (anonymous):

yeah. even if it's x inside. look:\[\frac{ d }{ dx }\sqrt{x}=\frac{ d }{ dx }x^{\frac{ 1 }{ 2 }}=x^{-\frac{ 1 }{ 2 }}\cdot(1)=x^{-\frac{ 1 }{ 2 }}\]

OpenStudy (yttrium):

@savi56 , try to take a look at this reference. There are some problems here that uses chain rule of differentiation. Try to analyze. This may help you. :) Just try the basics first. I think that will make you understand the rule better.

OpenStudy (anonymous):

after a while, you'll learn to recognize

OpenStudy (anonymous):

for example a function like f(z)=\[\sqrt{z}\] is too simple to use the chain rule and power rule would be better instead

OpenStudy (anonymous):

dear pgpilot, thank you! dear Yttrium, got it in a new tab thank you!

OpenStudy (anonymous):

take a look at this http://tutorial.math.lamar.edu/Classes/CalcI/ChainRule.aspx

OpenStudy (anonymous):

dear azbits, ahhhh i see because its simple. of course if it was the square root of x^2 the answer would be x. if it was the square root of x^3 then you would have to use the chain rule right?

OpenStudy (anonymous):

you can use the chain rule on any function... \[\frac{ d }{ dx }x=\frac{ d }{ dx }x^1=1\cdot x^0 \cdot(1) = 1\]

OpenStudy (anonymous):

i see! ok i get it now thank you guys for all your help! ill be back like next week for another question lol!

OpenStudy (anonymous):

YA'YA'LL ARE SMART lol. thank you all so much!

OpenStudy (anonymous):

well you could muliply the exponents and use power rule but you can use chain rule as well. Point is there is more than one way but some functions require you to use chain rule like the one on that website I linked.

OpenStudy (anonymous):

@savi56 if it was (2x-3) inside or (2x^3-3x^2+5) inside. if it's a pure power, rewrite using exponents and just use power rule.

OpenStudy (anonymous):

dear azbits, alrighty thats true dear pgpiolet326, ok because of the big exponents! hahahaha calculus i get you now!

OpenStudy (anonymous):

keep at it!

OpenStudy (anonymous):

i will lol thanks @azbits and @pgpilot326 !

OpenStudy (anonymous):

just practice with more problems and you'll figure it out :).

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

yeah im redoing my homework LMAO! night guys thanks!

OpenStudy (anonymous):

@azbits i like your profile picture btw :3 Daft Punk!

OpenStudy (anonymous):

thanks savi!

OpenStudy (anonymous):

np! :)

OpenStudy (anonymous):

were up all night to get calculus :D

OpenStudy (anonymous):

HAHA i like that xD! were up all night to defeat calculus is how i wud say it :P

OpenStudy (anonymous):

wud the derivative of 2x-(x^2)(tanx) by using the chain rule be 2-(2x^3)(tanx)(sec^2)(x)

OpenStudy (anonymous):

i just wanna make sure im right

OpenStudy (anonymous):

no you have a product in the 2nd term so you'd use the product rule for that.